# Semi-Bandit Learning for Monotone Stochastic Optimization

### v1

@@ 2 The Online Framework
$\alpha\text{-}R(T)=\sum_{t=1}^{T}\mathbb{E}[R^{t}(\mathbf{h}^{t-1})]\leq knf_{max}\sqrt{6\log(knT)}\,+\,Tf_{max}\cdot\mathbf{P}[\overline{G}]\leq knf_{max}\sqrt{6\log(knT)}+f_{max}.$
This completes the proof of Theorem 1.1 for up-monotone problems.
@@ 2.2 Proving the Stability Lemma
We now prove Lemma 2.3. Recall that there are two product distributions $\mathbf{E}=\{\mathcal{E}_{i}\}_{i=1}^{n}$ and $\mathbf{D}=\{\mathcal{D}_{i}\}_{i=1}^{n}$ with $\mathbf{D}\preceq_{\mathsf{SD}}\mathbf{E}$ and $\mathtt{TV}(\mathcal{E}_{i},\mathcal{D}_{i})\leq\epsilon_{i}$ for each $i\in[n]$. Let $\sigma=\mathtt{ALG}(\mathbf{E})$ be the policy returned by the (offline) $\alpha$-approximation algorithm, and $\sigma^{*}$ be an optimal policy under distribution $\mathbf{D}$. We want to upper bound $f(\sigma)-\alpha\cdot f(\sigma^{*})$. Let $\mathcal{C}$ denote the set of all feasible policies to the given instance. Then,
@@
Inequality (10) uses the fact that $\sigma$ is an $\alpha$-approximate policy to the instance with distribution $\mathbf{E}$. In (11), the equality uses the fact that $\sigma^{*}$ is an optimal policy for the instance with distribution $\mathbf{D}$, and the inequality is by up-monotonicity (Definition 1.1) and $\mathbf{D}\preceq_{\mathsf{SD}}\mathbf{E}$.
@@
Based on the above view of $\sigma$, we use $\mathcal{D}_{v}$ (resp. $\mathcal{E}_{v}$) to denote the independent distribution at each node $v\in\sigma$ under the joint distribution $\mathbf{D}$ (resp. $\mathbf{E}$). Using the above indexing of the nodes in $\sigma$, we define the following hybrid product distributions: for each $v\in[N]$ let $\mathbf{{H}}^{v}=\mathcal{D}_{1}\times\cdots\times\mathcal{D}_{v}\times\mathcal{E}_{v+1}\times\cdots\times\mathcal{E}_{N}$. Observe that $\mathbf{{H}}^{N}=\mathbf{{D}}$ and $\mathbf{{H}}^{0}=\mathbf{{E}}$. Using a telescoping sum, we can write:
@@
where $Q_{v}(\sigma)$ is the probability that $\sigma$ reaches node $v$ under distribution $\mathbf{D}$ and $\epsilon_{v}=\mathtt{TV}(\mathcal{D}_{v},\mathcal{E}_{v})$. We first complete the proof of the lemma using (12). Note that for any node $v$ labeled by item $i\in[n]$, we have $\epsilon_{v}=\mathtt{TV}(\mathcal{D}_{i},\mathcal{E}_{i})\leq\epsilon_{i}$.
@@
Towards proving (12), we introduce the following notation. Let $Q_{v}(\sigma\mid\mathbf{{U}})$ denote the probability that $\sigma$ reaches node $v$ under any distribution $\mathbf{{U}}$. Note that the event “$\sigma$ reaches node $v$” only depends on the random realizations at the ancestor nodes of $v$, which are all contained in $\{u\in[N]:u<v\}$ (by our indexing of nodes). Hence,
@@
[…]
We now bound the difference in expectation conditional on event ${\cal R}_{v}$. Let $\rho$ denote the subtree of $\sigma$ rooted at $v$ (including $v$ itself). We also use $\rho$ to denote the set of nodes in this subtree. Let $L$ denote all leaf nodes in subtree $\rho$, and for each $\ell\in L$ let $f_{\ell}$ be the function value accrued at leaf node $\ell$. For any realization $\{x_{w}:w\in\rho\}$, we use ($\ell(\{x_{w}:w\in\rho\})$ to denote the (unique) leaf that is reached when subtree $\rho$ is executed under this realization: this corresponds to following the branch labeled $x_{w}$ out of each node $w\in\rho$ (starting from node $v$). We now define function $h$ that maps any realization $\{x_{w}:w\in\rho\}$ to the function value $f_{k}$ at the leaf $k=\ell(\{x_{w}:w\in\rho\})$. By our assumption on the objective function of the stochastic problem, $h$ is bounded between 00 and $f_{max}$. We now define two product distributions on nodes of $\rho$:
@@
$h(\mathbf{{V}}_{1})=\mathbb{E}_{\mathbf{x}\sim\mathbf{{H}}^{v}}\left[f(\sigma,\mathbf{x})\mid{\cal R}_{v}\right]\quad\mbox{ and }\quad h(\mathbf{{V}}_{2})=\mathbb{E}_{\mathbf{x}\sim\mathbf{{H}}^{v-1}}\left[f(\sigma,\mathbf{x})\mid{\cal R}_{v}\right]$
@@
$\bigg{|}\mathbb{E}_{\mathbf{x}\sim\mathbf{{H}}^{v}}\left[f(\sigma,\mathbf{x})\mid{\cal R}_{v}\right]-\mathbb{E}_{\mathbf{x}\sim\mathbf{{H}}^{v-1}}\left[f(\sigma,\mathbf{x})\mid{\cal R}_{v}\right]\bigg{|}\,\,\leq\,\,f_{max}\cdot\mathtt{TV}(\mathbf{{V}}_{1},\mathbf{{V}}_{2})\,\leq\,f_{max}\cdot\mathtt{TV}(\mathcal{D}_{v},\mathcal{E}_{v}).$ (16)
@@
$\leq Q_{v}(\sigma)\cdot f_{max}\cdot\mathtt{TV}(\mathcal{D}_{v},\mathcal{E}_{v})=f_{max}\cdot Q_{v}(\sigma)\cdot\epsilon_{v}.$
@@ 2.3 Calculating a Stochastically Dominating Empirical Distribution
In order to motivate this algorithm, consider the simple case that $\mathcal{D}$ is a Bernoulli distribution, i.e., $\mathcal{D}=\mathtt{Bern}(p)$ where $p$ is the (unknown) probability of realizing to value $1$. We first compute the empirical mean $\widehat{p}$ of $p$ from the $m$ i.i.d. samples. Then, we set $\mathcal{E}=\mathtt{Bern}\left(\widehat{p}+\sqrt{\frac{\log(1/\delta)}{m}}\right)$, where $\delta>0$ is the algorithm’s failure probability. It is easy to see, by Chernoff-Hoeffding bounds, that $|p-\widehat{p}|<\sqrt{\frac{\log(1/\delta)}{m}}$ and $\mathcal{D}\preceq_{\mathsf{SD}}\mathcal{E}$ with probability at least $1-\delta$. Indeed, this is the strategy used in many algorithms that use the principle of optimism under uncertainty, such as the upper-confidence-bound (UCB) algorithm [LS20]. We will prove Theorem 2.1 by giving such a construction for any distribution $\mathcal{D}$ with finite support. We note that similar constructions have been used before, e.g. [GHZ19], but their notion of approximating $\mathbf{D}$ is different from our $\mathtt{TV}$-distance condition.
Consider any distribution $\mathcal{D}$ with support $\{a_{1},\cdots,a_{k}\}$ where $a_{1}<\cdots<a_{k}$, and let $\{p_{i}\}_{i=1}^{k}$ denote the respective probabilities and $\mathsf{X}$ the corresponding r.v. That is, $\mathbf{P}(\mathsf{X}=a_{i})=p_{i}$ for all $i\in[k]$. The algorithm takes as input $m$ i.i.d. samples $x_{1},\ldots,x_{m}\sim\mathcal{D}$. It defines the empirical distribution as $\widehat{p}_{i}:=\frac{1}{m}\cdot\sum_{j=1}^{m}\mathbf{I}[\mathsf{X}_{j}=a_{i}]$ for all $i\in[k]$. Since this empirical distribution might not stochastically dominate $\mathcal{D}$, the algorithm shifts some mass from lower to higher values in the support. However, this shift of mass needs to ensure that the resulting distribution is close to $\mathcal{D}$ in terms of total variation ($\mathtt{TV}$) distance. Algorithm 2 describes the entire procedure.
@@
2:let $\epsilon=\sqrt{\frac{\log(2k/\delta)}{2m}}$
3:for $i\in[k]$, let $\widehat{p}_{i}\leftarrow\frac{1}{m}\cdot\sum_{j=1}^{m}\mathbf{I}[\mathsf{X}_{j}=a_{i}]$
@@
We will now show that the empirical distribution $\mathcal{E}$ returned by our algorithm stochastically dominates $\mathcal{D}$ and is also a good approximation in terms of $\mathtt{TV}$ distance.
@@
We will show that $\bar{p}_{i}\geq 0$ for all $i\in[k]$ and $\sum_{i=1}^{k}\bar{p}_{i}=1$. If $\bar{p}_{k}=1$ then it is easy to see that $\mathcal{E}$ is a valid: all other probabilities $\bar{p}_{i}=0$. If $\bar{p}_{k}<1$ then $\widehat{p}_{k}<1-\epsilon$, which implies $\sum_{i=1}^{k-1}\widehat{p}_{i}=1-\widehat{p}_{k}>\epsilon$. So, index $y\in[k-1]$, and
@@
For any $i\in[k]$, let $q_{i}=\sum_{j\geq i}p_{j}$ and $\widehat{q}_{i}=\sum_{j\geq i}\widehat{p}_{j}$. Using Hoeffding’s bound and $\epsilon=\sqrt{\frac{\log(2k/\delta)}{2m}}$,
$\mathbf{P}\left(|\widehat{q}_{i}-q_{i}|\geq\epsilon\right)\leq 2e^{-2m\epsilon^{2}}\leq\frac{\delta}{k},\qquad\forall i\in[k].$
Using the union bound we have,
$\mathbf{P}\left(\exists i\in[k]\,:\,|\widehat{q}_{i}-q_{i}|\geq\epsilon\right)\leq\delta.$
Henceforth, we assume the “good” event that $|\widehat{q}_{i}-q_{i}|<\epsilon$ for all $i\in[k]$, which occurs with probability at least $1-\delta$. Under this event, we will show that $\mathcal{D}\preceq_{\mathsf{SD}}\mathcal{E}$ and $\mathtt{TV}(\mathcal{E},\mathcal{D})<2k\epsilon$. This would complete the proof of Theorem 2.1.
@@
$\mathtt{TV}(\widehat{\mathcal{E}},\mathcal{D})<(k-1)\epsilon.$ (18)
@@
$\mathtt{TV}(\widehat{\mathcal{E}},\mathcal{D})=\frac{1}{2}\cdot\sum_{i=1}^{k}|\widehat{p}_{i}-p_{i}|<(k-1)\epsilon,$
@@
    $\mathtt{TV}(\widehat{\mathcal{E}},\mathcal{E})=\frac{1}{2}\cdot\sum_{i=1}^{k}|\bar{p}_{i}-\widehat{p}_{i}|=\frac{1}{2}\left(|1-\widehat{p}_{k}|+\sum_{i=1}^{k-1}\widehat{p}_{i}\right)\leq\epsilon.$
@@
    $\mathtt{TV}(\widehat{\mathcal{E}},\mathcal{E})=\frac{1}{2}\cdot\sum_{i=1}^{k}|\bar{p}_{i}-\widehat{p}_{i}|=\frac{1}{2}\left(\epsilon+|\sum_{i=1}^{y}\widehat{p}_{i}-\epsilon-\widehat{p}_{y}|+\sum_{i=1}^{y-1}\widehat{p}_{i}\right)=\epsilon,$
@@
Hence, in either case, we have $\mathtt{TV}(\widehat{\mathcal{E}},\mathcal{E})\leq\epsilon$, which combined with (18) implies
$\mathtt{TV}(\mathcal{E},\mathcal{D})\leq\mathtt{TV}(\widehat{\mathcal{E}},\mathcal{E})+\mathtt{TV}(\widehat{\mathcal{E}},\mathcal{D})<k\epsilon.$
@@
    $\mathbf{P}(\mathsf{Z}\geq v)=1\geq\mathbf{P}(\mathsf{X}\geq v),$
    where $\mathsf{X}$ and $\mathsf{Z}$ have distributions $\mathcal{D}$ and $\mathcal{E}$ respectively. If $v>a_{y}$, then let $i\geq y+1$ be the smallest index such that $a_{i}\geq v$. We have
    $\mathbf{P}(\mathsf{Z}\geq v)=\sum_{j=i}^{k}\bar{p}_{j}=\sum_{j=i}^{k-1}\widehat{p}_{j}+\widehat{p}_{k}+\epsilon=\widehat{q}_{i}+\epsilon\geq q_{i}=\mathbf{P}(\mathsf{X}\geq v)\,,$

### v2

@@ 2 The Online Framework
$\alpha\text{-}R(T)=\sum_{t=1}^{T}\mathbb{E}[R^{t}(\mathbf{h}^{t-1})]\leq knf_{max}\sqrt{6T\log(nT)}\,+\,Tf_{max}\cdot\Pr[\overline{G}]\leq knf_{max}\sqrt{6T\log(nT)}+f_{max}.$
The last inequality uses Lemma 2.2. This completes the proof of Theorem 1.1. ∎
@@ 2.1 Proving the Stability Lemma
We now prove Lemma 2.3. Recall that there are two product distributions $\mathbf{E}=\{\mathcal{E}_{i}\}_{i=1}^{n}$ and $\mathbf{D}=\{\mathcal{D}_{i}\}_{i=1}^{n}$ with $\mathbf{D}\preceq_{\mathsf{SD}}\mathbf{E}$ and $\operatorname{\mathtt{TV}-\mathtt{Dist}}(\mathcal{E}_{i},\mathcal{D}_{i})\leq\epsilon_{i}$ for each $i\in[n]$. Let $\sigma=\mathtt{ALG}(\mathbf{E})$ be the policy returned by the (offline) $\alpha$-approximation algorithm, and $\sigma^{*}$ be an optimal policy under distribution $\mathbf{D}$. We want to upper bound $f(\sigma)-\alpha\cdot f(\sigma^{*})$. Let $\mathcal{C}$ denote the set of all feasible policies to the given instance. Then,
@@
Inequality (10) uses the fact that $\sigma$ is an $\alpha$-approximate policy to the instance with distribution $\mathbf{E}$. In (11), the equality uses the fact that $\sigma^{*}$ is an optimal policy for the instance with distribution $\mathbf{D}$, and the final inequality is by up-monotonicity (Definition 1.1) and $\mathbf{D}\preceq_{\mathsf{SD}}\mathbf{E}$.
@@
Based on the above view of $\sigma$, we use $\mathcal{D}_{v}$ (resp. $\mathcal{E}_{v}$) to denote the independent distribution at each node $v\in\sigma$ under the joint distribution $\mathbf{D}$ (resp. $\mathbf{E}$). Using the above indexing of the nodes in $\sigma$, we define the following hybrid product distributions: for each $v\in[N]$ let $\mathbf{{H}}^{v}=\mathcal{D}_{1}\times\cdots\times\mathcal{D}_{v}\times\mathcal{E}_{v+1}\times\cdots\times\mathcal{E}_{N}$. Observe that $f(\sigma\mid\mathbf{{H}}^{N})=f(\sigma\mid\mathbf{{D}})$ and $f(\sigma\mid\mathbf{{H}}^{0})=f(\sigma\mid\mathbf{{E}})$. Using a telescoping sum, we can write:
@@
where $Q_{v}(\sigma)$ is the probability that $\sigma$ reaches node $v$ under distribution $\mathbf{D}$ and $\epsilon_{v}=\operatorname{\mathtt{TV}-\mathtt{Dist}}(\mathcal{D}_{v},\mathcal{E}_{v})$. We first complete the proof of the lemma using (12). Note that for any node $v$ labeled by item $i\in[n]$, we have $\epsilon_{v}=\operatorname{\mathtt{TV}-\mathtt{Dist}}(\mathcal{D}_{i},\mathcal{E}_{i})\leq\epsilon_{i}$.
@@
Towards proving (12), we introduce the following notation. Let $Q_{v}(\sigma\mid\mathbf{{U}})$ denote the probability that $\sigma$ reaches node $v$ under distribution $\mathbf{{U}}$. Note that the event $\sigma$ reaches node vv” only depends on the random realizations at the ancestor nodes of $v$, which are all contained in $\{u\in[N]:u<v\}$ (by our indexing of nodes). Hence,
@@
$\mathbb{E}_{\mathbf{x}\sim\mathbf{{H}}^{v}}\left[f(\sigma,\mathbf{x})\mid\neg{\cal R}_{v}\right]-\mathbb{E}_{\mathbf{x}\sim\mathbf{{H}}^{v-1}}\left[f(\sigma,\mathbf{x})\mid\neg{\cal R}_{v}\right]\,\,=\,\,0.$ (15)
We now bound the difference in expectation conditional on event ${\cal R}_{v}$. Let $\rho$ denote the subtree of $\sigma$ rooted at $v$ (including $v$ itself). We also use $\rho$ to denote the set of nodes in this subtree. Let $L$ denote all leaf nodes in subtree $\rho$, and for each $\ell\in L$ let $f_{\ell}$ be the function value accrued at leaf node $\ell$. For any realization $\{x_{w}:w\in\rho\}$, we use $\ell(\{x_{w}:w\in\rho\})$ to denote the (unique) leaf that is reached when subtree $\rho$ is executed under this realization. This corresponds to following the branch labeled $x_{w}$ out of each node $w\in\rho$ (starting from node $v$). We now define function $h$ that maps any realization $\{x_{w}:w\in\rho\}$ to the function value $f_{k}$ at the leaf $k=\ell(\{x_{w}:w\in\rho\})$. By our assumption on the objective function of the stochastic problem, $h$ is bounded between 0 and $f_{max}$. We now define two product distributions on nodes of $\rho$:
@@
[…]
@@
$\bigg{|}\mathbb{E}_{\mathbf{x}\sim\mathbf{{H}}^{v}}\left[f(\sigma,\mathbf{x})\mid{\cal R}_{v}\right]-\mathbb{E}_{\mathbf{x}\sim\mathbf{{H}}^{v-1}}\left[f(\sigma,\mathbf{x})\mid{\cal R}_{v}\right]\bigg{|}\,\,$ $\leq\,\,f_{max}\cdot\operatorname{\mathtt{TV}-\mathtt{Dist}}(\mathbf{{V}}_{1},\mathbf{{V}}_{2})$
$\leq\,\,f_{max}\cdot\operatorname{\mathtt{TV}-\mathtt{Dist}}(\mathcal{D}_{v},\mathcal{E}_{v}).$ (16)
@@
$\leq Q_{v}(\sigma)\cdot f_{max}\cdot\operatorname{\mathtt{TV}-\mathtt{Dist}}(\mathcal{D}_{v},\mathcal{E}_{v})=f_{max}\cdot Q_{v}(\sigma)\cdot\epsilon_{v}.$
@@ 2.2 Calculating a Stochastically Dominating Empirical Distribution
In order to motivate this algorithm, consider the simple case that $\mathcal{D}$ is a Bernoulli distribution, i.e., $\mathcal{D}=\mathtt{Bern}(p)$ where $p$ is the (unknown) probability of realizing to value $1$. We first compute the empirical mean $\widehat{p}$ of $p$ from the $m$ i.i.d. samples. Then, we set $\mathcal{E}=\mathtt{Bern}\left(\widehat{p}+\sqrt{\frac{\log(1/\delta)}{m}}\right)$, where $\delta>0$ is the algorithm’s failure probability. It is easy to see, by Chernoff-Hoeffding bounds, that $|p-\widehat{p}|<\sqrt{\frac{\log(1/\delta)}{m}}$ and $\mathcal{D}\preceq_{\mathsf{SD}}\mathcal{E}$ with probability at least $1-\delta$. Indeed, this is the strategy used in many algorithms that use the principle of optimism under uncertainty, such as the upper-confidence-bound (UCB) algorithm (Auer et al., 2002). We will prove Theorem 2.1 by giving such a construction for any distribution $\mathcal{D}$ with finite support. We note that similar constructions have been used before, for e.g. in Guo et al. (2019), but their notion of approximating $\mathbf{D}$ is different from our $\operatorname{\mathtt{TV}-\mathtt{Dist}}$ condition.
Consider any distribution $\mathcal{D}$ with support $\{a_{1},\cdots,a_{k}\}$ where $a_{1}<\cdots<a_{k}$, and let $\{p_{i}\}_{i=1}^{k}$ denote the respective probabilities and $\mathsf{X}$ the corresponding r.v. That is, $\Pr(\mathsf{X}=a_{i})=p_{i}$ for all $i\in[k]$. The algorithm takes as input $m$ i.i.d. samples $x_{1},\ldots,x_{m}\sim\mathcal{D}$. It defines the empirical distribution as $\widehat{p}_{i}:=\frac{1}{m}\cdot\sum_{j=1}^{m}\mathbf{I}[x_{j}=a_{i}]$ for all $i\in[k]$. Since this empirical distribution might not stochastically dominate $\mathcal{D}$, the algorithm shifts some mass from lower to higher values in the support. However, this shift of mass needs to ensure that the resulting distribution is close to $\mathcal{D}$ in terms of $\operatorname{\mathtt{TV}-\mathtt{Dist}}$. Algorithm 2 describes the entire procedure.
@@
2:let $\epsilon=\sqrt{\frac{\log(2/\delta)}{2m}}$
3:for $i\in[k]$, let $\widehat{p}_{i}\leftarrow\frac{1}{m}\cdot\sum_{j=1}^{m}\mathbf{I}[x_{j}=a_{i}]$
@@
We will now show that the distribution $\mathcal{E}$ returned by our algorithm stochastically dominates $\mathcal{D}$ and is also a good approximation in terms of $\operatorname{\mathtt{TV}-\mathtt{Dist}}$.
@@
We will show that $\bar{p}_{i}\geq 0$ for all $i\in[k]$ and $\sum_{i=1}^{k}\bar{p}_{i}=1$. If $\bar{p}_{k}=1$ then it is easy to see that $\mathcal{E}$ is valid: all other probabilities $\bar{p}_{i}=0$. If $\bar{p}_{k}<1$ then $\widehat{p}_{k}<1-\epsilon$, which implies $\sum_{i=1}^{k-1}\widehat{p}_{i}=1-\widehat{p}_{k}>\epsilon$. So, index $y\in[k-1]$, and
@@
For any $i\in[k]$, let $q_{i}=\sum_{j\geq i}p_{j}$ and $\widehat{q}_{i}=\sum_{j\geq i}\widehat{p}_{j}$. Note that $\{q_{i}\}$ and $\{\widehat{q}_{i}\}$ denote the tail cumulative-distribution-functions (CDFs) of $\mathcal{D}$ and the empirical distribution respectively. Using the DKW inequality (Theorem 1.4), we obtain
$\Pr\left(\exists i\in[k]\,:\,|\widehat{q}_{i}-q_{i}|\geq\epsilon\right)\leq\delta,$
where $\epsilon=\sqrt{\frac{\log(2/\delta)}{2m}}$. Henceforth, we assume the “good” event that $|\widehat{q}_{i}-q_{i}|<\epsilon$ for all $i\in[k]$, which occurs with probability at least $1-\delta$. Under this event, we will show that $\mathcal{D}\preceq_{\mathsf{SD}}\mathcal{E}$ and $\operatorname{\mathtt{TV}-\mathtt{Dist}}(\mathcal{E},\mathcal{D})<k\epsilon$. This would complete the proof of Theorem 2.1.
@@
$\operatorname{\mathtt{TV}-\mathtt{Dist}}(\widehat{\mathcal{E}},\mathcal{D})<(k-1)\epsilon.$ (18)
@@
$\operatorname{\mathtt{TV}-\mathtt{Dist}}(\widehat{\mathcal{E}},\mathcal{D})\,\,=\,\,\frac{1}{2}\cdot\sum_{i=1}^{k}|\widehat{p}_{i}-p_{i}|\,\,<\,\,(k-1)\epsilon,$
@@
    $\operatorname{\mathtt{TV}-\mathtt{Dist}}(\widehat{\mathcal{E}},\mathcal{E})\,\,=\,\,\frac{1}{2}\cdot\sum_{i=1}^{k}|\bar{p}_{i}-\widehat{p}_{i}|\,\,=\,\,\frac{1}{2}\left(|1-\widehat{p}_{k}|+\sum_{i=1}^{k-1}\widehat{p}_{i}\right)\,\,\leq\,\,\epsilon.$
@@
    $\operatorname{\mathtt{TV}-\mathtt{Dist}}(\widehat{\mathcal{E}},\mathcal{E})=\frac{1}{2}\cdot\sum_{i=1}^{k}|\bar{p}_{i}-\widehat{p}_{i}|\,\,=\,\,\frac{1}{2}\left(\epsilon+|\sum_{i=1}^{y}\widehat{p}_{i}-\epsilon-\widehat{p}_{y}|+\sum_{i=1}^{y-1}\widehat{p}_{i}\right)=\epsilon,$
@@
Hence, in either case, we have $\operatorname{\mathtt{TV}-\mathtt{Dist}}(\widehat{\mathcal{E}},\mathcal{E})\leq\epsilon$, which combined with (18) implies
$\operatorname{\mathtt{TV}-\mathtt{Dist}}(\mathcal{E},\mathcal{D})\,\,\leq\,\,\operatorname{\mathtt{TV}-\mathtt{Dist}}(\widehat{\mathcal{E}},\mathcal{E})+\operatorname{\mathtt{TV}-\mathtt{Dist}}(\widehat{\mathcal{E}},\mathcal{D})\,\,<\,\,k\epsilon.$
@@
    $\Pr(\mathsf{Z}\geq v)=1\geq\Pr(\mathsf{X}\geq v),$
    where $\mathsf{X}\sim\mathcal{D}$ and $\mathsf{Z}\sim\mathcal{E}$. If $v>a_{y}$, then let $i\geq y+1$ be the smallest index such that $a_{i}\geq v$. We have
    $\Pr(\mathsf{Z}\geq v)=\sum_{j=i}^{k}\bar{p}_{j}=\sum_{j=i}^{k-1}\widehat{p}_{j}+\widehat{p}_{k}+\epsilon=\widehat{q}_{i}+\epsilon\geq q_{i}=\Pr(\mathsf{X}\geq v)\,$,